OpenStudy (anonymous):

x-intercept=2; containing the point (-5,4), using either the general form or the slope-intercept form of the equation of a line.

OpenStudy (anonymous):

x intercept 2 means you know that \[(2,0)\] is on the graph

OpenStudy (anonymous):

so you have two points, \[(2,0),(-5,4)\] find the slope via \[m=\frac{y_2-y_1}{x_2-x_1}\] and then use the point slope formula \[y-y_1=m(x-x_1)\]

OpenStudy (anonymous):

clear or you need it worked out?

OpenStudy (anonymous):

wait? my bad its (4,-5)

OpenStudy (anonymous):

ok fine, then the two points are \[(2,0)\] and \[4,-5)\] but the formulas are still the same

OpenStudy (anonymous):

\[m=\frac{0-(-5)}{2-4}=\frac{5}{-2}=-\frac{5}{2}\]

OpenStudy (anonymous):

is the answer y=-5/2-5?

OpenStudy (anonymous):

close

OpenStudy (anonymous):

+5?

OpenStudy (anonymous):

it is \[y=-\frac{5}{2}x+5\]

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

\[y-0=-\frac{5}{2}(x-2)\] \[y=-\frac{5}{2}x+5\] right, that is correct

OpenStudy (anonymous):

can you help me with this problem?

OpenStudy (anonymous):

The endpoints of the diameter of a circle are (-3,2) and (5,-6). Find the center and radius of the circle. Write the general equation of this circle.

OpenStudy (anonymous):

name it

OpenStudy (anonymous):

sure no problem

OpenStudy (anonymous):

you know it should look like \[(x-h)^2+(y-k)^2=r^2\] where the center is (h,k) and the radius is r

OpenStudy (anonymous):

so what we need is the center and the square of the radius. the center is the midpoint of the line, which you find by adding each coordinate and dividing by 2 (take the average)

OpenStudy (anonymous):

endpoints are (-3,2) and (5,-6). so the center is \[(\frac{-3+5}{2},\frac{2-6}{2})\]or \[(1,-2)\]

OpenStudy (anonymous):

(1,-2)

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

so now we know it looks like \[(x-1)^2+(y+2)^2=r^2\] and we only need \[r^2\]

OpenStudy (anonymous):

the square of the distance between (1,-2) and (-3,2) is \[(1+3)^2+(2+2)^2=4^2+4^2=16+16=32\]

OpenStudy (anonymous):

so your "final answer" is \[(x-1)^2+(y+2)^2=32\]

OpenStudy (anonymous):

you did the distance formula right?